In the future, parents will no longer need to push
Answer #1

Data:

m=40\, kg

l=3\, m

x=30\, cm=0.3\, m

\theta =45^{\circ}

Part a)

Applying the conservation of energy we have:

E=0

or

E_{i}=E_{f}

U_{ei}=U_{g}+K_{f}

The speed of the kid at the highest point (45 degrees) is zero, so; K_{f}=0

U_{ei}=U_{g}

\frac{1}{2}kx^{2}=mgH

Isolating k:

k=\frac{2mgH}{x^{2}}\, \, \, \, \, \, \, \, \, \, (1)

Where, H is the max height reaches for the kid, and is calulate using:

H =l-l\cos 45=3-3\cos 45

H =0.88\, m

Replacing values given in the statement of the problem and the value of H in equation (1) :

k=\frac{2mgH}{x^{2}}\, \, \, \, \, \, \, \, \, \, (1)

k=\frac{2(40\, kg)(9.8\, m/s^{2})(0.88\, m)}{(0.3\, m)^{2}}

k=7665.7\, N/m